OpenStudy (anonymous):

2x+7y=48 and 3x-12y=27

OpenStudy (texaschic101):

3x - 12y = 27 -- divide by 3 x - 4y = 9 x = 4y + 9 now you can sub that in for x in the other equation 2x + 7y = 48 2(4y + 9) + 7y = 48 8y + 18 + 7y = 48 8y + 7y = 48 - 18 15y = 30 y = 30/15 y = 2 now sub 2 in for y in either equation to find x x = 4y + 9 x = 4(2) + 9 x = 8 + 9 x = 17 check.. 3x - 12y = 27 3(17) - 12(2) = 27 51 - 24 = 27 27 = 27 (correct) solution : (17,2)